OpenStudy (anonymous):

Two angles are supplementary if their sum is 180 degrees. The larger angle measures six degrees more than twice the measure of a smaller angle. if x represents the measure of the smaller angle and these two angles are supplementary, find the measure of each angle.

OpenStudy (anonymous):



OpenStudy (anonymous):

isolate, substitute, solve

OpenStudy (anonymous):

substitute :) \[y=2x+6\]\[x+y=180\]put y into the second equation:\[x+(2x+6)=180\]add 2x and x:\[3x+6=180\]subtract 6 from 180:\[3x=174\]divide both sides by 3:\[x=~...?\] next whatever you get for x, put it into \(x+y=180\) and solve for y :)